# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v1

@@ I Introduction
Equally, QKMs have been studied from a theoretical perspective. Kübler et al. [25] provide insights into the inductive bias of quantum kernels and conclude that quantum speed-ups may only occur if one manages to encode knowledge about the problem at hand into underlying quantum circuits, while encoding the same bias into a classical model would be hard. In this regard, Refs.  [26, 27], the authors show that tuning the kernel’s bandwidth (which can be seen as a data preprocessing step) controls the model’s inductive bias. Consequently, these works identify the quantum kernel bandwidth as the key hyperparameter controlling the expressiveness of the model and give a theory, which shows that varying the bandwidth enables generalization. However, the exponential size of the quantum feature space can hinder generalization and cause exponentially concentrated quantum kernel values [17, 28]. To circumvent this setback, Huang et al. [16] introduced the family of PQKs. [^1] We note that Ref. [28] also derives exponential concentration bounds for PQKs. However, in practice PQKs usually suffer less from exponentially large Hilbert spaces, provided certain assumption on corresponding data encoding circuits are taken into account [17]. Their key findings include proving that classical learners can approximate quantum outputs with enough data and introducing methods for assessing potential quantum advantages. An illustrative application of these concepts is given in Ref. [29], where the authors provide numerical evidence against quantum advantage for FQKs on classical data. They show that tuning the kernel’s bandwidth improve the model performance and thus enables generalization, but also results in classically tractable kernels due to unfavorable geometric difference values.
@@
In this work, we conduct an extensive benchmarking study for classification and regression tasks using various QKMs such as quantum support vector classification (QSVC), quantum kernel ridge regression (QKRR), and quantum support vector regression (QSVR). Additionally, we generate a broad and general database by exploring nine popular data encoding circuits from the literature, offering a diverse analysis on the impact of the choice of encoding circuit. Beyond that, we propose two different encoding strategies to introduce feature redundancies and systematically compare them. Furthermore, the present work delves deep into the different design aspects of PQKs by examining the effects of measurement operators and outer kernels on model performance. We additionally address the question of which components of the PQK definition are responsible for learning, i.e., the projected quantum circuits versus the outer kernel. Finally, we numerically support all findings by rigorous correlation analysis. In total, this study resulted in over $20,000$ quantum kernel models that were trained to provide an extensive database and to ensure robust and comprehensive insights. To facilitate this, we developed a software tool. Its basic functionality is shown in Fig. 1.
@@
Previous studies have focused mostly on classification tasks. For example, Bowles et al. [34] conducted a large-scale evaluation of popular QML models, including Quantum Neural Networks (QNNs), FQKs and PQKs across various datasets. While QKMs where part of the study, the influence of the different design choices was not a particular focus. Their work was preceded by first attempts to systematically explore certain aspects of quantum model design, cf., e.g., Refs [35, 36]. Egginger et al. [37] have conducted a hyperparameter study for QKMs which extends the findings of Ref. [29] for PQKs. However, unlike this work, they exclusively focus on a single feature map (Hamiltonian evolution). In contrast to these previous findings, the present study offers an in-depth analysis of QKMs that significantly broadens the scope of investigation in multiple aspects and aims to answers aspects that have been previously untouched. This includes addressing regression problems, encoding mechanisms and a detailed analyis of PQKs. By extending previous findings we aim to contribute to a holistic understanding of QKMs.
@@
This work is organized as follows. Section II introduces the theoretical basics of QKMs. Hereafter, we give detailed insights about the study design, including an overview on the multitude of different quantum kernel models considered in this work, an introduction of the datasets as well as implementation details for the corresponding hyperparameter optimization pipeline and the final experimental setup. In Sec. IV we first thoroughly investigate hyperparameter importances and model performances for classification and regression tasks across all datasets, encoding circuits and quantum kernel models and support corresponding findings with correlation analyses. Secondly, we carefully examine different data encoding strategies and give an in-depth analysis on different design options within the PQK approach, i.e., choice of outer kernel function and measurement operator. We discuss universal findings and patterns across all experiments and comment on the necessity of entanglement in data encoding circuits in Sec. V.
@@ II Theoretical Background
One of the most interesting aspects of QKMs is that they can be formally embedded into the rich and powerful mathematical framework of classical kernel theory [3, 11]. The key idea behind the conventional kernelized approach to (supervised) machine learning is to find and analyze patterns by transforming the respective learning problem from the original input data domain $\mathcal{X}$ to a higher-dimensional (potentially infinite-dimensional) feature space $\mathcal{F}$, where the learning tasks attains a trivial form, i.e., becomes a linear model. This mapping is accomplished by a feature map $\phi:\mathcal{X}\to\mathcal{F};\mathbf{x}\mapsto\phi(\mathbf{x})$. Kernels, are real- or complex-valued symmetric and positive semi-definite functions of two input data points, $k:\mathcal{X}\times\mathcal{X}\to\mathbb{K}$, where $\mathbb{K}$ can be $\mathbb{R}$ or $\mathbb{C}$, respectively. In this regard, another central concept is that of reproducing kernel Hilbert space (RKHS) $\mathcal{H}_{k}$, which is an alternative feature spaces of a kernel; more precisely a feature space of functions constructed from the kernel. The RKHS uniquely determines the kernel and vice versa [38] and in addition, for every kernel there exists at least one feature map such that [11]
@@
To formalize this, we define the density matrix $\rho_{\bm{\theta}}(\mathbf{x})=\outerproduct{\psi(\mathbf{x},\bm{\theta})}{\psi(\mathbf{x},\bm{\theta})}$ | as the corresponding data-encoding feature map [11, 4]. With this, we can leverages the native geometry of the quantum state space to defining a quantum kernel in a natural way using the Hilbert-Schmidt inner product, i.e.,
@@
where $\gamma\in\mathbb{R_{+}}$ is a hyperparameter and $\rho_{\bm{\theta}}^{k}(\mathbf{x})$ is the 1-RDM, i.e., the partial trace of the quantum state $\rho_{\bm{\theta}}(\mathbf{x})$ over all qubits except for the $k$-th qubit; cf. Eq. (6). Here, we introduced the latter notation for later use in Sec. IV.4. Note that unless otherwise stated, we are referring to the form as given in Eq. (9) when considering PQKs.
@@ III.1 Models
  • Regularization parameters, these are the Tikhonov regularization strengths $\lambda$ and $C$ for QKRR and QSVR/QSVC, respectively. Additionally, for QSVR there is an additional hyperparameter $\varepsilon$, which specifies the range within which no penalty is associated in the training loss function with a point predicted within a distance $\varepsilon$ from the actual value
@@
We point out that data encoding unitaries (cf. Eq. (3)) additionally have multiple degrees of freedom. Data encoding circuits are at the heart of each QML method and significantly influence core properties of the resulting model. Generating problem-specific encoding circuits with, e.g., proper gate sets and corresponding structure marks a distinct research branch, cf., e.g., Refs. [43, 24]. This is beyond the scope of this work, wherefore we restrict ourselves to nine data encoding circuits from the QML literature [44, 5, 7, 45, 31, 27, 28, 46]. In this context, we randomly initialize the variationally trainable parameters of the encoding circuits that involve $\bm{\theta}$ (cf. Eq. (3)) using a fixed seed. For details on these encoding circuit as well as corresponding illustrations, we refer to Appendix B.
@@ III.2 Datasets
Selecting meaningful datasets for the sake of a conclusive study constitutes a highly nontrivial task. Therefore, our choice is mainly driven by aiming for datasets that are not too easy and whose complexity is ideally adjustable. Additionally, the size and dimensionality needs to be suited for QML applications with reasonable simulation time. Unless stated otherwise we used $M=240$ training- and $M^{\prime}=60$ test data points within each dataset.
For studying classification, we use two binary classification datasets [^2] The resulting classification problems are balanced. introduced in Refs. [47, 48]. These datasets have been also used a previous benchmarking study [34].
@@
This dataset describes the curvature and distance of two one-dimensional curves embedded into a $d$-dimensional space. We follow the data generation procedure of Ref. [34] and use low-degree ($D$) Fourier series to embed two sets of data sampled from a 1-d interval as curves into $d$ dimensions while adding Gaussian noise $\sigma=0.01$. The respective complexity can be controlled by fixing $d=4$ and vary $D=2,\ldots,20$, while adapting the offset $\Delta=\nicefrac{{1}}{{2D}}$.
@@
where $x_{i}$ with $i\in\{1,2,3,4,5\}$ is the $i$-th component of the data point $\mathbf{x}$ and $\sigma$ is the standard deviation of Gaussian noise applied to the output. We set $\sigma=0.01$ for all simulations. Only $n=5$ features contribute to the computation of the labels, while the remaining features are independent. Due to its non-linearity as well as its interaction between features and noise characteristics, this dataset mimics real-world regression scenarios. For this study, we generate datasets for $d=5,\ldots 15$ features.
@@
This dataset is based on the fashion-MNIST dataset [53], which is sent through a quantum circuit to create a quantum-based dataset. We use the procedure as described in Ref. [8]. For this, first the fashion-MNIST feature dimension is reduced by principle component analysis. The resulting $d$ features are then encoded into a quantum state using an $n$-qubit encoding circuit as proposed by Havlicek et al. [4]. The subsequent arbitrary single qubit rotations as applied in Ref. [8] are omitted. Finally, the labels for are generated by computing expectation values w.r.t. measuring the first qubit in the Pauli-$Z$ basis. We generate datasets with $d=2,\ldots 15$ principal components.
@@
This real world dataset provides taken from Ref. [54], which is a vast benchmark database for machine learning potential energy surfaces of a wide range of different molecules. Here, we use the ammonia data labeled with “STATIC-g32n1-1M” and “DZERO”. While the original dataset is in $xyz$-format we transformed them into internal coordinates (i.e. bond lengths and angles). Since nonlinear molecules with $N$ atoms show $f=3N-6$ degrees of freedom, this dataset consists of six features with corresponding ground state energies (in Hartee) computed at CCSD(F12*)(T)/cc-pVDZ-F12 level of theory defining the labels. The dataset consists of 193 samples from which $M=155$ are used for training and $M^{\prime}=38$ for testing.
For the Friedman#1 and QFMNIST regression problems the number of features can be viewed as control parameter to adjust the dataset complexity (cf. Fig. 3).
@@ III.3 Experimental setup and Implementation
In the following we summarize the experimental setup for the hyperparameter search and provide some insights into the respective implementation.
All simulations in this study are based on sQUlearn with the PennyLane [55] statevector simulator device.
@@
We use sQUlearn for evaluating quantum kernel Gram matrices, which provides FQKs according to Eq. (5) and allows for defining PQKs as generally given in Eq. (8). For PQKs, we investigate the impact of different outer kernel functions $\kappa$. Specifically, we consider the Gaussian (RBF) kernel (cf., Eq. (9), which is the default in sQUlearn), the Matérn kernel [41] with $\nu=1.5$, i.e.,
@@
Moreover, we study the effect of using different $k$-local measurement operators for defining PQKs (cf., Eqs. (II) and (8)). We use the 1-RDM on all qubits w.r.t. different (combinations of) Pauli operators
@@
The QKMs as implemented in sQUlearn work analogously to their classical counterparts in scikit-learn. In this work we use QKRR and QSVR for regression tasks and QSVC for solving classification problems. The corresponding regularization hyperparameters are $\lambda$, $C$ and $\varepsilon$ as well as $C$, respectively.
@@
We develop the tool OKMTuner to facilitate the extensive hyperparameter search of this study. The code is based on sQUlearn [31], optuna [32] and scikit-learn [33]. QKMTuner consists of two main routines: a hyperparameter optimization within a grid search and a method for studying hyperparameter importances; cf. Fig. 1. The grid-search is given a list of data encoding circuits for each of which it builds a user-defined $n_{\mathrm{qubits}}\times n_{\mathrm{layers}}$-grid on which a hyperparameter search is performed for each grid point and each data encoding circuit for a given dataset. For evaluating hyperparameter importances, the respective method only takes a list of data encoding circuits and automatically searches for the optimal hyperparameters including $n_{\mathrm{qubits}}$ and $n_{\mathrm{layers}}$. In both cases, the hyperparameters are determined by maximizing the minimum between the mean and median of five-fold cross-validation scores. This choice of the objective function prevents the scores from becoming too optimistic and thus helps to prevent overfitting towards an easy fold. The respective scoring method can be set manually, whereas in this work we use the area under the receiver operating characteristic curve (ROC-AUC) from prediction scores for classification and (negative) mean squared error (MSE) for regression tasks, respectively. For hyperparameter sampling we use the tree-structured parzen estimator [57], albeit different algorithms as provided by optuna are supported. The code of QKMTuner is available in Ref. [30].

### v2

@@ I Introduction
Equally, QKMs have been studied from a theoretical perspective. Kübler et al. [25] provide insights into the inductive bias of quantum kernels and conclude that quantum speed-ups may only occur if one manages to encode knowledge about the problem at hand into underlying quantum circuits, while encoding the same bias into a classical model would be hard. In this regard, in Refs. [26, 27], the authors show that tuning the kernel’s bandwidth (which can be seen as a data preprocessing step) controls the model’s inductive bias. Consequently, these works identify the quantum kernel bandwidth as the key hyperparameter controlling the expressiveness of the model and give a theory, which shows that varying the bandwidth enables generalization. However, the exponential size of the quantum feature space can hinder generalization and cause exponentially concentrated quantum kernel values [17, 28]. To circumvent this setback, Huang et al. [16] introduced the family of PQKs. [^1] We note that Ref. [28] also derives exponential concentration bounds for PQKs. However, in practice PQKs usually suffer less from exponentially large Hilbert spaces, provided certain assumptions on corresponding data encoding circuits are taken into account [17]. Their key findings include proving that classical learners can approximate quantum outputs with enough data and introducing methods for assessing potential quantum advantages. An illustrative application of these concepts is given in Ref. [29], where the authors provide numerical evidence against quantum advantage for FQKs on classical data. They show that tuning the kernel’s bandwidth improve the model performance and thus enables generalization, but also results in classically tractable kernels due to unfavorable geometric difference values.
@@
We conduct an extensive benchmarking study for classification and regression tasks using various QKMs such as quantum support vector classification (QSVC), quantum kernel ridge regression (QKRR), and quantum support vector regression (QSVR). Additionally, we generate a broad and general database by exploring nine popular data encoding circuits from the literature, offering a diverse analysis on the impact of the choice of encoding circuit. Beyond that, we propose two different encoding strategies to introduce feature redundancies and systematically compare them. Furthermore, the present work delves deep into the different design aspects of PQKs by examining the effects of measurement operators and outer kernels on model performance. We additionally address the question of which components of the PQK definition are responsible for learning, i.e., the projected quantum circuits versus the outer kernel. Finally, we numerically support all findings by rigorous correlation analysis. In total, this study resulted in over $20,000$ quantum kernel models that were trained to provide an extensive database and to ensure robust and comprehensive insights. To facilitate this, we developed a software tool. Its basic functionality is indicated in Fig. 1.
@@
Previous studies have focused mostly on classification tasks. For example, Bowles et al. [34] conducted a large-scale evaluation of popular QML models, including Quantum Neural Networks (QNNs), FQKs and PQKs across various datasets. While QKMs where part of the study, the influence of the different design choices was not a particular focus. Their work was preceded by first attempts to systematically explore certain aspects of quantum model design, cf., e.g., Refs [35, 36]. Egginger et al. [37] have conducted a hyperparameter study for QKMs which extends the findings of Ref. [29] for PQKs. However, unlike this work, they exclusively focus on a single feature map (Hamiltonian evolution). In contrast to these previous findings, the present study offers an in-depth analysis of QKMs that significantly broadens the scope of investigation in multiple aspects and aims to answer aspects that have been previously untouched. This includes addressing regression problems, encoding mechanisms and a detailed analysis of PQKs. By extending previous findings we aim to contribute to a holistic understanding of QKMs.
@@
This work is organized as follows. Section II introduces the theoretical basics of QKMs. Hereafter, we give detailed insights about the study design, including an overview on the multitude of different quantum kernel models considered in this work, an introduction of the datasets as well as implementation details for the corresponding hyperparameter optimization pipeline and the final experimental setup. In Sec. IV we first thoroughly investigate model performances and the influence of hyperparameters for classification and regression tasks across all datasets, encoding circuits and quantum kernel models and support corresponding findings with correlation analyses. Secondly, we carefully examine different data encoding strategies and give an in-depth analysis on different design options within the PQK approach, i.e., choice of outer kernel function and measurement operator. We discuss universal findings and patterns across all experiments and comment on the necessity of entanglement in data encoding circuits in Sec. V.
@@ II Theoretical Background
One of the most interesting aspects of QKMs is that they can be formally embedded into the rich and powerful mathematical framework of classical kernel theory [3, 11]. The key idea behind the conventional kernelized approach to (supervised) machine learning is to find and analyze patterns by transforming the respective learning problem from the original input data domain $\mathcal{X}$ to a higher-dimensional (potentially infinite-dimensional) feature space $\mathcal{F}$, where the learning tasks attains a trivial form, i.e., becomes a linear model. This mapping is accomplished by a feature map $\phi:\mathcal{X}\to\mathcal{F};\mathbf{x}\mapsto\phi(\mathbf{x})$. Kernels, are real- or complex-valued symmetric and positive semi-definite functions of two input data points, $k:\mathcal{X}\times\mathcal{X}\to\mathbb{K}$, where $\mathbb{K}$ can be $\mathbb{R}$ or $\mathbb{C}$, respectively. In this regard, another central concept is that of reproducing kernel Hilbert space (RKHS) $\mathcal{H}_{k}$, which is an alternative feature space of a kernel; more precisely a feature space of functions constructed from the kernel. The RKHS uniquely determines the kernel and vice versa [38] and in addition, for every kernel there exists at least one feature map such that [11]
@@
To formalize this, we define the density matrix $\rho_{\bm{\theta}}(\mathbf{x})=\outerproduct{\psi(\mathbf{x},\bm{\theta})}{\psi(\mathbf{x},\bm{\theta})}$ | as the corresponding data-encoding feature map [11, 4]. With this, we can leverages the native geometry of the quantum state space and naturally define a quantum kernel using the Hilbert-Schmidt inner product, i.e.,
@@
where $\gamma\in\mathbb{R_{+}}$ is a hyperparameter and $\rho_{\bm{\theta}}^{k}(\mathbf{x})$ is the 1-RDM, i.e., the partial trace of the quantum state $\rho_{\bm{\theta}}(\mathbf{x})$ over all qubits except for the $k$-th qubit; cf. Eq. (6). Here, we introduced the latter notation for later use in Sec. IV.5. Note that unless otherwise stated, we are referring to the form as given in Eq. (9) when considering PQKs.
@@ III.1 Models
  • Regularization parameters, these are the Tikhonov regularization strengths $\lambda$ and $C$ for QKRR and QSVR/QSVC, respectively. Additionally, for QSVR there is also the hyperparameter $\varepsilon$, which specifies the range within which no penalty is associated in the training loss function with a point predicted within a distance $\varepsilon$ from the actual value
@@
We point out that data encoding unitaries (cf. Eq. (3)) additionally have multiple degrees of freedom. Data encoding circuits are at the heart of each QML method and significantly influence core properties of the resulting model. Generating problem-specific encoding circuits with, e.g., proper gate sets and corresponding structure marks a distinct research branch, cf., e.g., Refs. [43, 24]. This is beyond the scope of this work, wherefore we restrict ourselves to nine data encoding circuits from the QML literature [44, 5, 7, 45, 31, 27, 28, 46]. In this context, we randomly initialize the variationally trainable parameters of those encoding circuits that involve $\bm{\theta}$ (cf. Eq. (3)) using a fixed seed. For details on these encoding circuits as well as corresponding illustrations, we refer to Appendix B.
@@ III.2 Datasets
Selecting meaningful datasets for the sake of a conclusive study constitutes a highly nontrivial task. Therefore, our choice is mainly driven by aiming for datasets that are not too easy and whose complexity is ideally adjustable. Additionally, the size and dimensionality needs to be suited for QML applications with reasonable simulation time. Unless stated otherwise, we use $M=240$ training- and $M^{\prime}=60$ test data points within each dataset.
For studying classification, we use two binary classification datasets [^2] The resulting classification problems are balanced. introduced in Refs. [47, 48]. These datasets have been also used in a previous benchmarking study [34].
@@
This dataset describes the curvature and distance of two one-dimensional curves embedded into a $d$-dimensional space. We follow the data generation procedure of Ref. [34] and use low-degree ($D$) Fourier series to embed two sets of data sampled from a 1-d interval as curves into $d$ dimensions, while adding Gaussian noise $\sigma=0.01$. The respective complexity can be controlled by fixing $d=4$ and vary $D=2,\ldots,20$, while adapting the offset $\Delta=\nicefrac{{1}}{{2D}}$.
@@
where $x_{i}$ with $i\in\{1,2,3,4,5\}$ is the $i$-th component of the data point $\mathbf{x}$ and $\sigma$ is the standard deviation of Gaussian noise applied to the output. We set $\sigma=0.01$ for all simulations. Only $d=5$ features contribute to the computation of the labels, while the remaining features are independent. Due to its non-linearity as well as its interaction between features and noise characteristics, this dataset mimics real-world regression scenarios. For this study, we generate datasets for $d=5,\ldots 15$ features.
@@
This dataset is based on the fashion-MNIST dataset [53], which is send through a quantum circuit to create a quantum-based dataset. We use the procedure as described in Ref. [8]. For this, first the fashion-MNIST feature dimension is reduced by principle component analysis. The resulting $d$ features are then encoded into a quantum state using an $d$-qubit encoding circuit as proposed by Havlicek et al. [4]. The subsequent arbitrary single qubit rotations as applied in Ref. [8] are omitted. Finally, the labels are generated by computing expectation values w.r.t. measuring the first qubit in the Pauli-$Z$ basis. We generate datasets with $d=2,\ldots 15$ principal components.
@@
This real world dataset is taken from Ref. [54], which is a vast benchmark database for machine learning potential energy surfaces of a wide range of different molecules. Here, we use the ammonia data labeled with “STATIC-g32n1-1M” and “DZERO”. While the original dataset is in $xyz$-format we transformed them into internal coordinates (i.e. bond lengths and angles). Since nonlinear molecules with $N$ atoms show $f=3N-6$ degrees of freedom, this dataset consists of $d=6$ features with corresponding ground state energies (in Hartee) computed at CCSD(F12*)(T)/cc-pVDZ-F12 level of theory defining the labels. The dataset consists of 193 samples from which $M=155$ are used for training and $M^{\prime}=38$ for testing.
For the Friedman#1 and QFMNIST regression problems the number of features $d$ can be viewed as control parameter to adjust the dataset complexity (cf. Fig. 3).
@@ III.3 Experimental setup and Implementation
In the following, we summarize the experimental setup for the hyperparameter search and provide some insights into the respective implementation.
All simulations in this study are based on sQUlearn [31] with the PennyLane [55] statevector simulator device.
@@
We use sQUlearn for evaluating quantum kernel Gram matrices, which provides FQKs according to Eq. (5) and allows for defining PQKs as generally given in Eq. (8). For PQKs, we investigate the impact of different outer kernel functions $\kappa$. Specifically, we consider the Gaussian (RBF) kernel (cf., Eq. (9), which is the default in sQUlearn), the Matérn kernel [41] with $\nu=\nicefrac{{3}}{{2}}$, i.e.,
@@
Moreover, we study the effect of using different $k$-local measurement operators $O^{k}$ for defining PQKs (cf., Eqs. (II) and (8)). We use the 1-RDM on all qubits w.r.t. different (combinations of) Pauli operators
@@
The QKMs as implemented in sQUlearn work analogously to their classical counterparts in scikit-learn. In this work, we use QKRR and QSVR for regression tasks and QSVC for solving classification problems. The corresponding regularization hyperparameters are $\lambda$, $C$ and $\varepsilon$ as well as $C$, respectively as introduced above.
@@
We develop the tool QKMTuner to facilitate the extensive hyperparameter search of this study. The code is based on sQUlearn [31], Optuna [32] and scikit-learn [33]. QKMTuner consists of two main routines: a hyperparameter optimization within a grid search and a method for studying hyperparameter importances; cf. Fig. 1. The grid-search is given a list of data encoding circuits for each of which it builds a user-defined $n_{\mathrm{qubits}}\times n_{\mathrm{layers}}$-grid on which a hyperparameter search is performed for each grid point and each data encoding circuit for a given dataset. For evaluating hyperparameter importances, the respective method only takes a list of data encoding circuits and automatically searches for the optimal hyperparameters including $n_{\mathrm{qubits}}$ and $n_{\mathrm{layers}}$. In both cases, the hyperparameters are determined by maximizing the minimum between the mean and median of five-fold cross-validation scores. This choice of the objective function prevents the scores from becoming too optimistic and thus helps to prevent overfitting towards an easy fold. The respective scoring method can be set manually, whereas in this work we use the area under the receiver operating characteristic curve (ROC-AUC) from prediction scores for classification and the (negative) mean squared error (MSE) for regression tasks, respectively. For hyperparameter sampling we use the tree-structured parzen estimator [57], albeit different algorithms as provided by Optuna are supported. The code of QKMTuner is available via GitLab [30].
@@
IV Results
In this section, we report all findings from the different types of experiments that have been performed, before we discuss these results in Sec. V.
IV.1 Model Performance
To start with, we delve into the relationships concerning the generalizability of QKMs. For studying this, we fix $n_{\mathrm{qubits}}$ for every dataset to the respective number of features for all upcoming investigations and perform hyperparameter optimizations for every data encoding circuit with $n_{\mathrm{layers}}\in[1,8]$
This analysis of test performances of different QKMs as a function of increasing dataset complexity is given in Fig. 4, which illustrates the results for the Friedman, QFMNIST, two curves diff, and hidden manifold diff datasets. In these cases, results for each QKM and dataset within the respective dataset family are aggregated across the data encoding circuits with corresponding obtained optimal $n_{\mathrm{layers}}^{*}$, yielding the best MSE or ROC-AUC score, respectively.
A comparison with Fig. 3 indicates that the behavior of test performance scores as a function of a control parameter qualitatively aligns with the increasing complexity of the respective datasets. This confirms that $\overline{C}$ is a useful measure of complexity. The comparison between FQK and PQK approaches in Figs. 4 (a) and (b), particularly for larger problem instances with increasing $d$, which transfers to increasing $n_{\mathrm{qubits}}$, reveals no significant performance differences. While PQKs demonstrate slightly superior performance, the difference is not as substantial as one might anticipate given the challenges posed in connection with exponential concentration [28], which might already be apparent as we approach 15 qubits. The absence of this observation may be explained by the presence of the bandwidth tuning parameter $w_{\mathrm{e}}$, which shows significant correlations with performance scores as discussed below. The comparison between QKRR and QSVR, each equipped with either a FQK or a PQK, does not reveal a clear advantage for either method. Notably, the QFMNIST ($d=2$) dataset constitutes the only case for which we observe a clear performance difference between FQK and PQK. Interestingly, this marks the most complex regression task, cf., Fig. 3. Whether this is merely a coincidence or actually due to a more fundamental mechanism might be interesting to investigate in the future. Regarding the classification tasks, we do not observe any significant difference between both QKMs.
We generally observe some outliers exhibiting particularly poor test performance scores. The comparison with associated training performance scores in Fig. A10 shows that this is probably a result of overfitting for certain model combinations. Since we investigate a large variety of combinations, it is unavoidable that some of the combinations result in models that are too expressive for some tasks. As shown in Appendix D.1 this is most dominant for QKRR-FQK/PQK approaches with increasing dataset complexity, while QSVR appears to be more robust. We point out here that our test size is quite small ($M^{\prime}=60$, as mentioned in Sec. III.2), which represents a clear limitation towards more detailed statements in this direction. This should be carefully considered in future studies.
Figure 4: Overview of test performance scores of respective QKMs as a function of increasing dataset complexity. Results within each dataset are aggregated across all data encoding circuits with corresponding optimal $n_{\mathrm{layers}}^{*}$ yielding minimum/maximum test performance scores for regression/classification, respectively. The upper panel displays two regression tasks, where the MSE is used to measure the prediction accuracy. The Friedman dataset family is shown in (a), where the number of features can be viewed as control parameter for dataset complexity. The QFMNIST dataset family is shown in (b), where the number of principal components used in the dataset generation procedure controls the dataset complexity. The lower panel illustrates the two classification tasks of this study, where we use the ROC-AUC score to assess classification accuracy. In (c) we show the two curves diff dataset family with the degree $D$ controlling the complexity. The hidden manifold diff family is given in (d) with the manifold dimension $m$ as respective control parameter.
IV.2 Influence of Hyperparameters